# Improved Sample Complexity For Diffusion Model Training Without Empirical Risk Minimizer Access

### v5

@@ Appendix A Comparison with Prior Works
In this section, we provide a detailed comparison of our results with prior work. Specifically, we analyze the sample complexity bounds presented in Gupta et al. (2024), and show how combining their results with those of Block et al. (2020) leads to an alternative bound.
@@
We begin by examining the sample complexity claim of $\mathcal{O}(1/\epsilon^{3})$ reported in Gupta et al. (2024). A closer analysis reveals that the actual sample complexity is $\widetilde{\mathcal{O}}(1/\epsilon^{5})$, once the error over all the discretization steps is properly accounted for and a union bound is applied.
The main result regarding the sample complexity of estimating the score function as given in Theorem C.2 of Gupta et al. (2024) is as follows. We re-iterate this Theorem here.
Theorem C.2 Gupta et al. (2024). Let $q$ be a distribution of $\mathbb{R}^{d}$ with second moment $m_{2}^{2}$. Let $\phi_{\theta}(\cdot)$ be the fully connected neural network with ReLU activations parameterized by $\theta$, with $P$ total parameters and depth $D$. Let $\Theta>1$. For any $\gamma>0$, there exist $K=\widetilde{\mathcal{O}}\left(\frac{d}{\epsilon^{2}+\delta^{2}}\log^{2}\frac{m_{2}+1/m_{2}}{\gamma}\right)$ discretization times $0=t_{0}<\cdots<t_{K}<T$ such that if for each $t_{k}$, there exists some score function $\hat{s}_{\theta}$ with $\|\theta^{*}\|_{\infty}\leq\Theta$ such that
@@
Note that Lemma B.6 of Gupta et al. (2024) states that
@@
In contrast, our analysis reduces the overall complexity by a factor of $\widetilde{\mathcal{O}}(1/\epsilon)$, yielding the tightest known bounds for neural score estimation in diffusion models, i.e., $\widetilde{\mathcal{O}}(1/\epsilon^{4})$. Further, unlike Gupta et al. (2024), our analysis avoids using the $1-\delta$-quantile bound on the score norm and instead directly bounds the global $L^{2}$ score estimation error thus avoids applying a union bound across time steps, and finally achieve tighter sample complexity guarantees.
@@ C.1 Bounding the Statistical Error
We now bound terms (I) and (II) using generalization results. From Lemma 5 (Theorem 26.5 of Shalev-Shwartz & Ben-David (2014)), if the loss function $\widehat{\mathcal{L}}(\theta)$ is uniformly bounded over the parameter space $\Theta^{\prime\prime}=\{\theta_{k}^{a},\theta_{k}^{b}\}$, then with probability at least $1-\delta$, we have
@@
Note that the functions $v_{t}(x)$ and $v_{\theta}(x,t)$ are uniformly bounded. Thus using Theorem 26.5 of Shalev-Shwartz & Ben-David (2014) we have with probability at least $1-\delta$,
@@ C.2 Bounding Optimization Error
Define the expected suboptimality $e_{i}:=\mathbb{E}\bigl[\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star}\bigr]$.
@@
$\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\|^{2}\mid\theta_{i}\bigr]=\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})-\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}\mid\theta_{i}\bigr]\leq\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2},$
@@
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}\right]\leq\mathcal{L}_{k}(\theta_{i})-\eta_{i}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\Bigl(\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2}\Bigr).$ (100)
@@
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}\right]\leq\mathcal{L}_{k}(\theta_{i})-\frac{\eta_{i}}{2}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\sigma^{2}.$ (101)
@@
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})-\mathcal{L}_{k}^{\star}\mid\theta_{i}\right]\leq\bigl(1-\mu\eta_{i}\bigr)\bigl(\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star}\bigr)+\frac{L}{2}\eta_{i}^{2}\sigma^{2}.$
@@ Appendix D Intermediate Lemmas
Lemma 5 (Theorem 26.5 of Shalev-Shwartz & Ben-David (2014)).
@@
Lemma 6 (Extewnsion of Massart’s Lemma Bousquet et al. (2003)).

### v6

@@ 4 Conclusion and Future Work
The Supplementary contains additional technical material, including full proofs of our main results, detailed explanations of key lemmas, algorithmic descriptions, and further discussion to support and extend the main text.
@@ Appendix A Comparison with Prior Works
In this section, we provide a detailed comparison of our results with prior work. Specifically, we analyze the sample complexity bounds presented in guptaimproved, and show how combining their results with those of block2020generative leads to an alternative bound.
@@
We begin by examining the sample complexity claim of $\mathcal{O}(1/\epsilon^{3})$ reported in guptaimproved. A closer analysis reveals that the actual sample complexity is $\widetilde{\mathcal{O}}(1/\epsilon^{5})$, once the error over all the discretization steps is properly accounted for and a union bound is applied.
The main result regarding the sample complexity of estimating the score function as given in Theorem C.2 of guptaimproved is as follows. We re-iterate this Theorem here.
Theorem C.2 guptaimproved. Let $q$ be a distribution of $\mathbb{R}^{d}$ with second moment $m_{2}^{2}$. Let $\phi_{\theta}(\cdot)$ be the fully connected neural network with ReLU activations parameterized by $\theta$, with $P$ total parameters and depth $D$. Let $\Theta>1$. For any $\gamma>0$, there exist $K=\widetilde{\mathcal{O}}\left(\frac{d}{\epsilon^{2}+\delta^{2}}\log^{2}\frac{m_{2}+1/m_{2}}{\gamma}\right)$ discretization times $0=t_{0}<\cdots<t_{K}<T$ such that if for each $t_{k}$, there exists some score function $\hat{s}_{\theta}$ with $\|\theta^{*}\|_{\infty}\leq\Theta$ such that
@@
Note that Lemma B.6 of guptaimproved states that
@@
In contrast, our analysis reduces the overall complexity by a factor of $\widetilde{\mathcal{O}}(1/\epsilon)$, yielding the tightest known bounds for neural score estimation in diffusion models, i.e., $\widetilde{\mathcal{O}}(1/\epsilon^{4})$. Further, unlike guptaimproved, our analysis avoids using the $1-\delta$-quantile bound on the score norm and instead directly bounds the global $L^{2}$ score estimation error thus avoids applying a union bound across time steps, and finally achieve tighter sample complexity guarantees.
@@ C.1 Bounding the Statistical Error
We now bound terms (I) and (II) using generalization results. From Lemma 5 (Theorem 26.5 of shalev2014understanding), if the loss function $\widehat{\mathcal{L}}(\theta)$ is uniformly bounded over the parameter space $\Theta^{\prime\prime}=\{\theta_{k}^{a},\theta_{k}^{b}\}$, then with probability at least $1-\delta$, we have
@@
Note that the functions $v_{t}(x)$ and $v_{\theta}(x,t)$ are uniformly bounded. Thus using Theorem 26.5 of shalev2014understanding we have with probability at least $1-\delta$,
@@ C.2 Bounding Optimization Error
Define the expected suboptimality $e_{i}:=\bigl[\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star}\bigr]$.
@@
$\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\|^{2}\bigr]=\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})-\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}\bigr]\leq\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2},$
@@
$\left[\mathcal{L}_{k}(\theta_{i+1})\right]\leq\mathcal{L}_{k}(\theta_{i})-\eta_{i}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\Bigl(\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2}\Bigr).$ (100)
@@
$\left[\mathcal{L}_{k}(\theta_{i+1})\right]\leq\mathcal{L}_{k}(\theta_{i})-\frac{\eta_{i}}{2}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\sigma^{2}.$ (101)
@@
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})-\mathcal{L}_{k}^{\star}\right]\leq\bigl(1-\mu\eta_{i}\bigr)\bigl(\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star}\bigr)+\frac{L}{2}\eta_{i}^{2}\sigma^{2}.$
@@ Appendix D Intermediate Lemmas
Lemma 5 (Theorem 26.5 of shalev2014understanding).
@@
Lemma 6 (Extewnsion of Massart’s Lemma bousquet2003introduction).
@@
SGD under PL with Heavy-Tailed Noise via Gradient Clipping
Setup.
We minimize a differentiable objective $f:\mathbb{R}^{d}\to\mathbb{R}$ with (possibly nonconvex) $\mu$–PL property:
$\frac{1}{2}\|\nabla f(x)\|^{2}\;\geq\;\mu\big(f(x)-f^{\star}\big)\quad\forall x,$ (125)
and LL–smoothness:
$f(y)\;\leq\;f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2}\quad\forall x,y.$ (126)
At iterate $x_{t}$ we observe a (conditionally) unbiased stochastic gradient $g_{t}=g(x_{t},\xi_{t})$:
$\mathbb{E}[g_{t}\mid x_{t}]\;=\;\nabla f(x_{t}),$ (127)
but without bounded variance. Instead we assume a finite $q$-th moment for some $q\in(1,2]$:
$\mathbb{E}\!\left[\|g_{t}-\nabla f(x_{t})\|^{q}\,\middle|\,x_{t}\right]\;\leq\;\sigma_{q}^{q}.$ (128)
Algorithm (clipped SGD).
Fix stepsizes $\{\eta_{t}\}_{t\geq 0}$ and clipping thresholds $\{\tau_{t}\}_{t\geq 0}$. Define the clipped gradient
$\tilde{g}_{t}\;=\;\mathrm{clip}(g_{t},\tau_{t})\;:=\;\begin{cases}g_{t},&\|g_{t}\|\leq\tau_{t},\\[3.0pt] \dfrac{\tau_{t}}{\|g_{t}\|}\,g_{t},&\|g_{t}\|>\tau_{t},\end{cases}$
and update $x_{t+1}=x_{t}-\eta_{t}\tilde{g}_{t}$.
We write $\Delta_{t}:=f(x_{t})-f^{\star}$ and take expectations over all randomness.
Assumption 6 (Choices of schedules).
Let $c\in(0,\,\frac{1}{2L}]$, $\gamma\geq 1$, and pick
$\eta_{t}\;=\;\frac{c}{t+\gamma},\qquad\tau_{t}\;=\;\Big(\frac{C_{1}}{L}\cdot\frac{t+\gamma}{\eta_{t}}\Big)^{\frac{1}{2q}}\sigma_{q}\;=\;\Big(\frac{C_{1}}{Lc}\Big)^{\frac{1}{2q}}\,\sigma_{q}\,(t+\gamma)^{\frac{1}{2q}},$ (129)
for a numerical constant $C_{1}\geq 1$ specified in the proof (any fixed choice suffices).
Main result. The theorem shows that (despite infinite variance) clipped SGD with the above schedules achieves the same $O(1/t)$ rate as in the bounded-variance case.
Theorem 3 (Heavy-tailed SGD under PL).
Suppose equation 125–equation 128 and Assumption 6 hold with $q\in(1,2]$. Then there exists a constant $K=K(L,\mu,c,\gamma,q,\sigma_{q})$ such that for all $t\geq 0$,
$\mathbb{E}[\Delta_{t}]\;\leq\;\frac{\max\{\mathbb{E}[\Delta_{0}],\,K\}}{t+\gamma}.$ (130)
Equivalently, by equation 125, $\mathbb{E}\|\nabla f(x_{t})\|^{2}=O(1/(t+\gamma))$.
Proof roadmap.
We first obtain a one-step descent inequality under clipping; we then bound the clipping bias using only the $q$-moment, choose schedules so that the additive noise terms decay strictly faster than $1/(t+\gamma)$, and solve the resulting recursion.
Lemma 9 (One-step progress with clipped gradients).
Under equation 126, for any $t$,
$\mathbb{E}[\Delta_{t+1}\mid x_{t}]\;\leq\;\Delta_{t}\;-\;\eta_{t}\langle\nabla f(x_{t}),\,\mathbb{E}[\tilde{g}_{t}\mid x_{t}]\rangle\;+\;\frac{L\eta_{t}^{2}}{2}\,\mathbb{E}\big[\|\tilde{g}_{t}\|^{2}\mid x_{t}\big].$ (131)
Moreover $\|\tilde{g}_{t}\|\leq\tau_{t}$, hence $\mathbb{E}[\|\tilde{g}_{t}\|^{2}\mid x_{t}]\leq\tau_{t}^{2}$.
Proof.
By $L$-smoothness with $y=x_{t+1}=x_{t}-\eta_{t}\tilde{g}_{t}$,
$f(x_{t+1})\leq f(x_{t})-\eta_{t}\langle\nabla f(x_{t}),\tilde{g}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\tilde{g}_{t}\|^{2}.$
Taking $\mathbb{E}[\cdot\mid x_{t}]$ gives equation 131. The norm bound follows from the definition of $\mathrm{clip}(\cdot,\tau_{t})$. ∎
Lemma 10 (Clipping bias from a $q$-moment).
Let $q\in(1,2]$ and suppose equation 127–equation 128 hold. Then there exists a universal constant $C_{q}\leq 2$ such that
$\big\|\mathbb{E}[\tilde{g}_{t}\mid x_{t}]-\nabla f(x_{t})\big\|\;\leq\;C_{q}\,\frac{\sigma_{q}^{q}}{\tau_{t}^{\,q-1}}.$ (132)
Proof.
Write $\delta_{t}:=g_{t}-\tilde{g}_{t}$. Then $\mathbb{E}[\tilde{g}_{t}\mid x_{t}]-\nabla f(x_{t})=\mathbb{E}[g_{t}-\delta_{t}\mid x_{t}]-\nabla f(x_{t})=-\mathbb{E}[\delta_{t}\mid x_{t}]$ by unbiasedness. Hence
$\big\|\mathbb{E}[\tilde{g}_{t}\mid x_{t}]-\nabla f(x_{t})\big\|\;=\;\big\|\mathbb{E}[\delta_{t}\mid x_{t}]\big\|\;\leq\;\mathbb{E}[\|\delta_{t}\|\mid x_{t}].$
If $\|g_{t}\|\leq\tau_{t}$, then $\delta_{t}=0$; otherwise $\tilde{g}_{t}=\tau_{t}g_{t}/\|g_{t}\|$ so that $\|\delta_{t}\|=\|g_{t}-\tilde{g}_{t}\|=\|g_{t}\|-\tau_{t}\leq\|g_{t}\|$. Therefore $\|\delta_{t}\|\leq\|g_{t}\|{1}\{\|g_{t}\|>\tau_{t}\}$ and
$\mathbb{E}[\|\delta_{t}\|\mid x_{t}]\;\leq\;\mathbb{E}\big[\|g_{t}\|\,{1}\{\|g_{t}\|>\tau_{t}\}\,\big|\,x_{t}\big]\;\leq\;\frac{\mathbb{E}[\|g_{t}\|^{q}\mid x_{t}]}{\tau_{t}^{\,q-1}}\;\leq\;\frac{\mathbb{E}[\|g_{t}-\nabla f(x_{t})\|^{q}\mid x_{t}]+\|\nabla f(x_{t})\|^{q}}{\tau_{t}^{\,q-1}}.$
Using $(a+b)^{q}\leq 2^{q-1}(a^{q}+b^{q})$ in reverse yields $\mathbb{E}[\|g_{t}\|^{q}\mid x_{t}]\leq 2^{q-1}\mathbb{E}[\|g_{t}-\nabla f(x_{t})\|^{q}\mid x_{t}]+2^{q-1}\|\nabla f(x_{t})\|^{q}$. Absorbing the harmless gradient term via $\|\nabla f(x_{t})\|^{q}\leq\|\nabla f(x_{t})\|^{2}+1$ (or, more simply, by enlarging the constant) together with equation 128 gives equation 132 with a constant $C_{q}\leq 2$. ∎
Lemma 11 (Descent recursion).
Under the previous assumptions, there exist constants
$A\;=\;\frac{C_{q}^{2}}{2},\qquad B\;=\;\frac{L}{2}$
such that
$\mathbb{E}[\Delta_{t+1}]\;\leq\;(1-\mu\eta_{t})\,\mathbb{E}[\Delta_{t}]\;+\;A\,\eta_{t}\,\frac{\sigma_{q}^{2q}}{\tau_{t}^{2(q-1)}}\;+\;B\,\eta_{t}^{2}\,\tau_{t}^{2}.$ (133)
Proof.
From Lemma 9 and Cauchy–Schwarz,
$\mathbb{E}[\Delta_{t+1}\mid x_{t}]\leq\Delta_{t}-\eta_{t}\langle\nabla f(x_{t}),\mathbb{E}[\tilde{g}_{t}\mid x_{t}]\rangle+\frac{L\eta_{t}^{2}}{2}\tau_{t}^{2}.$
Add and subtract $\nabla f(x_{t})$ inside the inner product:
$\eta_{t}\langle\nabla f(x_{t}),\mathbb{E}[\tilde{g}_{t}\mid x_{t}]\rangle\;=\;-\eta_{t}\|\nabla f(x_{t})\|^{2}+\eta_{t}\langle\nabla f(x_{t}),\nabla f(x_{t})-\mathbb{E}[\tilde{g}_{t}\mid x_{t}]\rangle.$
Bound the mixed term by Young’s inequality $ab\leq\tfrac{1}{2}a^{2}+\tfrac{1}{2}b^{2}$:
$\eta_{t}\langle\nabla f,\nabla f-\mathbb{E}[\tilde{g}_{t}]\rangle\;\leq\;\frac{\eta_{t}}{2}\|\nabla f(x_{t})\|^{2}+\frac{\eta_{t}}{2}\big\|\mathbb{E}[\tilde{g}_{t}\mid x_{t}]-\nabla f(x_{t})\big\|^{2}.$
Combining yields
$\mathbb{E}[\Delta_{t+1}\mid x_{t}]\leq\Delta_{t}-\frac{\eta_{t}}{2}\|\nabla f(x_{t})\|^{2}+\frac{\eta_{t}}{2}\big\|\mathbb{E}[\tilde{g}_{t}\mid x_{t}]-\nabla f(x_{t})\big\|^{2}+\frac{L\eta_{t}^{2}}{2}\tau_{t}^{2}.$
By PL, $\|\nabla f(x_{t})\|^{2}\geq 2\mu\Delta_{t}$; by Lemma 10, $\|\mathbb{E}[\tilde{g}_{t}\mid x_{t}]-\nabla f(x_{t})\|^{2}\leq C_{q}^{2}\sigma_{q}^{2q}\tau_{t}^{-2(q-1)}$. Taking unconditional expectation gives equation 133 with $A=C_{q}^{2}/2$, $B=L/2$. ∎
Lemma 12 (Balancing schedules).
With the schedules equation 129 there exists a constant $K_{0}$ such that for all $t$,
$A\,\eta_{t}\,\frac{\sigma_{q}^{2q}}{\tau_{t}^{2(q-1)}}\;+\;B\,\eta_{t}^{2}\,\tau_{t}^{2}\;\leq\;\frac{K_{0}}{(t+\gamma)^{\,2-\frac{1}{q}}}.$ (134)
Proof.
Let us choose $\tau_{t}$ by equating the two additive terms in equation 133:
$A\,\eta_{t}\,\frac{\sigma_{q}^{2q}}{\tau_{t}^{2(q-1)}}\;\approx\;B\,\eta_{t}^{2}\,\tau_{t}^{2}.$
This gives $\tau_{t}^{2q}=\frac{A}{B}\,\frac{\sigma_{q}^{2q}}{\eta_{t}}$ and hence $\tau_{t}=\big(\tfrac{A}{B}\big)^{\frac{1}{2q}}\sigma_{q}\,\eta_{t}^{-\frac{1}{2q}}$. Since $\eta_{t}=c/(t+\gamma)$, we obtain $\tau_{t}=\big(\tfrac{A}{Bc}\big)^{\frac{1}{2q}}\sigma_{q}(t+\gamma)^{\frac{1}{2q}}$, which is of the form equation 129 with $C_{1}=A$ (up to harmless constants). Substituting back, each additive term scales as
$A\,\eta_{t}\,\frac{\sigma_{q}^{2q}}{\tau_{t}^{2(q-1)}}\;\asymp\;(t+\gamma)^{-1}\,(t+\gamma)^{-\frac{q-1}{q}}\;=\;(t+\gamma)^{-\big(2-\frac{1}{q}\big)},$
and similarly for $B\,\eta_{t}^{2}\,\tau_{t}^{2}$. Absorbing constants into $K_{0}$ yields equation 134. ∎
Lemma 13 (A Robbins–Siegmund style recursion).
Let $\{s_{t}\}$ satisfy for some $\alpha>0$, $\gamma\geq 1$, $r>1$, and $u_{t}\geq 0$,
$s_{t+1}\;\leq\;\Big(1-\frac{\alpha}{t+\gamma}\Big)s_{t}\;+\;\frac{u_{t}}{(t+\gamma)^{r}},$
with $\sup_{t}u_{t}\leq U<\infty$. Then there exists $K_{\alpha}$ such that
$s_{t}\;\leq\;\frac{\max\{s_{0},\,K_{\alpha}U\}}{t+\gamma}\qquad\forall t.$
Proof.
This is standard; one can verify by induction using the comparison sequence $\bar{s}_{t}:=\frac{C}{t+\gamma}$ with $C$ chosen large enough so that
$\bar{s}_{t+1}-\Big(1-\frac{\alpha}{t+\gamma}\Big)\bar{s}_{t}\;=\;\frac{C}{t+1+\gamma}-\frac{C}{t+\gamma}+\frac{\alpha C}{(t+\gamma)^{2}}\;\geq\;\frac{C(\alpha-1)}{(t+\gamma)^{2}},$
while the perturbation is $O((t+\gamma)^{-r})$ with $r>1$; choose $C$ so the RHS dominates for all $t$. ∎
Proof of Theorem 3.
Combine Lemma 11 with Lemma 12 to obtain
$\mathbb{E}[\Delta_{t+1}]\;\leq\;\Big(1-\frac{\mu c}{t+\gamma}\Big)\mathbb{E}[\Delta_{t}]\;+\;\frac{K_{0}}{(t+\gamma)^{\,2-\frac{1}{q}}}.$
Since $q>1$, we have $2-\frac{1}{q}>1$, and Lemma 13 with $\alpha=\mu c$, $r=2-\frac{1}{q}$ yields $\mathbb{E}[\Delta_{t}]\leq\frac{\max\{\mathbb{E}[\Delta_{0}],K\}}{t+\gamma}$ for some $K$ depending only on the problem parameters. This proves equation 130. ∎
Remarks.
(i) The choice $\eta_{t}=\Theta(1/t)$ is critical for the PL contraction; the clipping threshold grows as $\tau_{t}=\Theta\!\big(\sigma_{q}\,t^{1/(2q)}\big)$, which balances the bias term $O\!\big(\eta_{t}\sigma_{q}^{2q}\tau_{t}^{-2(q-1)}\big)$ and the smoothness penalty $O(\eta_{t}^{2}\tau_{t}^{2})$ so that both decay at rate $(t+\gamma)^{-(2-1/q)}=o(1/t)$ when $q>1$.
(ii) For $q=2$ (finite variance), the analysis recovers the classical $O(1/t)$ rate with $\tau_{t}=\Theta(\sigma_{2}\,t^{1/4})$.
(iii) One can replace clipping by a robust mini-batch estimator (median-of-means, geometric median, or Catoni). Such estimators deliver sub-Gaussian deviations for the aggregated gradient under finite $q>1$ moments; plugging those into Lemma 9 with $\tau_{t}$ replaced by the aggregator’s confidence radius yields the same $O(1/t)$ rate with slightly different constants.
Theorem 4 (Last-iterate, expectation and a.s.).
Assume the setup from Theorem 3: $L$-smooth, $\mu$-PL, unbiased oracle, and a finite $q$-th moment for some $q\in(1,2]$. Run clipped SGD $x_{t+1}=x_{t}-\eta_{t}\tilde{g}_{t}$ with $\eta_{t}=c/(t+\gamma)$ and $\tau_{t}=\Theta\!\big(\sigma_{q}\,\eta_{t}^{-1/(2q)}\big)=\Theta\!\big(\sigma_{q}(t+\gamma)^{1/(2q)}\big)$.
Then:
(i) Expectation: $\qquad\mathbb{E}[f(x_{t})-f^{\star}]\;\leq\;\frac{C}{t+\gamma}.$ (135)
[…]
Moreover, $\sum_{t=0}^{\infty}\eta_{t}\,\mathbb{E}\|\nabla f(x_{t})\|^{2}<\infty$ and, a fortiori, $\mathbb{E}\|\nabla f(x_{t})\|^{2}=O(1/(t+\gamma))$.
Proof sketch.
(i) Expectation. The recursion proved earlier (Lemma 11)
$\mathbb{E}[\Delta_{t+1}]\leq(1-\mu\eta_{t})\mathbb{E}[\Delta_{t}]+A\,\eta_{t}\,\frac{\sigma_{q}^{2q}}{\tau_{t}^{2(q-1)}}+B\,\eta_{t}^{2}\tau_{t}^{2}$
already bounds the last iterate in expectation. With the balancing choice $\tau_{t}\propto\sigma_{q}\,\eta_{t}^{-1/(2q)}$ the additive terms decay like $(t+\gamma)^{-(2-1/q)}$ (Lemma 12), which is $o(1/(t+\gamma))$ for $q>1$. A Robbins–Siegmund style comparison (Lemma 13) gives equation 135.
(ii) Almost sure convergence. Decompose one step as
$\Delta_{t+1}\leq\Delta_{t}-\eta_{t}\langle\nabla f(x_{t}),\tilde{g}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\tilde{g}_{t}\|^{2}.$
Add and subtract $\mathbb{E}[\tilde{g}_{t}\mid\mathcal{F}_{t}]$ to expose a martingale increment:
$\eta_{t}\langle\nabla f(x_{t}),\tilde{g}_{t}\rangle=-\eta_{t}\|\nabla f(x_{t})\|^{2}+\eta_{t}\langle\nabla f(x_{t}),\nabla f(x_{t})-\mathbb{E}[\tilde{g}_{t}\mid\mathcal{F}_{t}]\rangle-\eta_{t}\langle\nabla f(x_{t}),M_{t+1}\rangle,$
where $M_{t+1}:=\tilde{g}_{t}-\mathbb{E}[\tilde{g}_{t}\mid\mathcal{F}_{t}]$ is a martingale difference. Using Young’s inequality and the bias bound (Lemma 10) yields
$\Delta_{t+1}\leq\Big(1-\mu\eta_{t}\Big)\Delta_{t}+a_{t}+b_{t}-\eta_{t}\langle\nabla f(x_{t}),M_{t+1}\rangle,$
with deterministic $a_{t}\asymp\eta_{t}\sigma_{q}^{2q}\tau_{t}^{-2(q-1)}$ and $b_{t}\asymp\eta_{t}^{2}\tau_{t}^{2}$. With our schedules, $\sum_{t}(a_{t}+b_{t})<\infty$ because $a_{t},b_{t}=\Theta((t+\gamma)^{-(2-1/q)})$ and 21/q>12-1/q>1 for $q>1$.
For the martingale term, $\|M_{t+1}\|\leq 2\tau_{t}$, so the quadratic variation is controlled by
$\sum_{t}\eta_{t}^{2}\mathbb{E}\!\left[\|M_{t+1}\|^{2}\mid\mathcal{F}_{t}\right]\;\lesssim\;\sum_{t}\eta_{t}^{2}\tau_{t}^{2}\;\asymp\;\sum_{t}(t+\gamma)^{-(2-1/q)}\;<\;\infty.$
Thus $\sum_{t}\eta_{t}\langle\nabla f(x_{t}),M_{t+1}\rangle$ converges a.s. (martingale convergence theorem), and the Robbins–Siegmund almost-supermartingale lemma applies, giving $\Delta_{t}\to\Delta_{\infty}$ a.s. and $\sum_{t}\eta_{t}\|\nabla f(x_{t})\|^{2}<\infty$ a.s.
Since $\sum_{t}\eta_{t}=\infty$ and $\eta_{t}>0$, the latter forces $\liminf_{t}\|\nabla f(x_{t})\|=0$; by PL, $\liminf_{t}\Delta_{t}=0$. But $\Delta_{t}$ is nonnegative and has a limit, so $\Delta_{\infty}=0$, i.e., $f(x_{t})\to f^{\star}$ a.s. PL then implies $x_{t}$ converges to the (global) stationary set. This proves equation 136. ∎
High-probability variant (sketch).
Because $\|\tilde{g}_{t}\|\leq\tau_{t}$ and $\sum_{t}\eta_{t}^{2}\tau_{t}^{2}<\infty$, Freedman/Azuma on the martingale $\sum_{s\leq t}\eta_{s}\langle\nabla f(x_{s}),M_{s+1}\rangle$ yields, for any $\delta\in(0,1)$,
$f(x_{t})-f^{\star}\;\leq\;\frac{C_{1}}{t+\gamma}\;+\;\frac{C_{2}\log(1/\delta)}{t+\gamma}\quad\text{with probability }\geq 1-\delta,$
for problem-dependent constants $C_{1},C_{2}$.